OpenStudy (anonymous):

How do you find the intersection of lines y1=-3/2x+3 and y2=-3/5x+2?

OpenStudy (anonymous):

Two lines intersect if they share a common point \((x,y)\). This means that when they intersect, both lines have the same x and y coordinates. Each y coordinate is expressed as a function of the x coordinate. At this intersection point, \(y_1=y_2\), so you simply solve the following equation for \(x\): \[-\frac{3}{2}x+3=-\frac{3}{5}x+2\] After you find this \(x\), you plug it into one of the given functions and solve for \(y\).

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

You're welcome!